# GL1-02-06-24-2-supplement

### v2

@@ 4.1 No Covariates
Suppose Assumption 1 holds and consider the model in (12). Also, suppose (unconditional) distributional parallel trends hold as in Definition 1. Then, $\eta_{gt}$ and $\lambda_{gt}$ are (point) identified if and only if $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are known.
Proposition 2 highlights the fact that in a setup with no further structure and no additional covariates, point identification of $\eta_{gt}$ and $\lambda_{gt}$ is impossible without the knowledge of the distribution function $F_{\varepsilon_{gt}}$. [^8] Unlike the if-and-only-if characterizations for distributional parallel trends in threshold-crossing models in this paper, Yamauchi (2020) establishes a point identification result for the model in (12) where $F_{\varepsilon_{gt}}$ is assumed to be known. The paper does, however, not consider the case where $F_{\varepsilon_{gt}}$ is unknown (see Subsection 4.2) or where the outcome is potentially misreported (see Section 5). The proposition can be straightforwardly generalized to conditional on covariates. We note that, with known $F_{\varepsilon_{gt}}$, the parameters are uniquely determined by the observable quantities
$\eta_{gt}=\frac{-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\quad\text{and}\quad\lambda_{gt}=\frac{1}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}$
for $(g,t)\neq(1,1)$. Moreover, Proposition 1 establishes that:
$\eta_{11}=\eta_{01}+\frac{\eta_{10}-\eta_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
Nevertheless, employing a parametric stand naturally calls for caution regarding potential misspecification. Therefore, we will ease this distributional assumption in Subsection 4.2 below.
4.2 Including Covariates
Next, we consider a semiparametric version of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and covariates $\mathbf{X}_{gt}$ enter the model. We will exploit variation in the latter to obtain a point identification result for the counterfactual $Y_{11}^{\ast}(0)$ distribution. At the same time, in order to facilitate the comparison with more standard ordered choice models from the literature and in light of the extension to misreporting in Section 5, we will place further structure on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively. Specifically, for every $(g,t)\in\{0,1\}$, we assume that, with probability one, $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}$. Thus:
@@ 4.2 Including Covariates
Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the median of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter.999 Note that interpretation of $\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ as conditional median will be important for the identification of the intercept parameter $\eta_{c,gt}$ (e.g., see Manski, 1988). Letting $\overline{\mathcal{X}}_{gt}$ the support of $\overline{\mathbf{X}}_{gt}$, we make the following assumptions:
Assumption 2.
For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\varepsilon_{gt}$ is independent of $\mathbf{X}_{gt}$.
Assumption 3.
For each $(g,t)\neq(1,1)$, there exists at least one $k\in\{1,\ldots$, $K\}$ such that $\eta_{k,gt}\neq 0$ and
$\operatorname{Pr}\left(X_{k,gt}\in\left(a_{1},a_{2}\right)\mid\overline{\mathbf{X}}_{-k,gt}\right)>0$
with probability one for all open intervals $\left(a_{1},a_{2}\right)\subseteq\mathbb{R}$, where $\overline{\mathbf{X}}_{-k,gt}$ denotes the subvector of $\overline{\mathbf{X}}_{gt}$ excluding the $k$th slope. W.l.o.g., let $k=1$.
Assumption 4.
For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\overline{\mathcal{X}}_{gt}$ is not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
Assumption 5.
It holds that $\mathcal{X}_{11}\subseteq\cap_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
Assumption 2 states covariate independence. Although this is often challenging to justify in practice, it remains a standard condition in the semiparametric estimation and identification literature, playing a crucial role in the identification argument (cf. Athey and Imbens, 2006). Assumption 3 on the other hand requires the existence of a “special regressor” with large support, an assumption frequently used in models with discrete responses to ensure point identification of the parameters (see Manski, 1988; Lewbel, 2000b; Han and Vytlacil, 2017; Khan et al., 2021). [^10] In principle, other identifying strategies from the semiparametric literature, such as exploiting the differentiability with respect to the continuous variables (e.g., Klein and Spady, 1993), are also possible with more complicated arguments. Finally, Assumption 4 is a standard identification assumption ruling out perfect multicollinearity, while Assumption 5 postulates a common support of $\mathbf{X}_{11}$ and the other subgroups.
We have the following identification result:
Proposition 3.
Suppose Assumptions 1-5 hold and consider the model in (13). Also, suppose (unconditional) distributional parallel trends hold as in Definition 1. Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$ and $F_{\varepsilon_{gt}}(v)$ are identified. In particular, it holds for the counterfactual median and scaling parameters that:
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
In contrasting the identification result of the counterfactual distribution in Proposition 3 with the point identification in Theorem 4.4 of Athey and Imbens (2006) for the discrete CiC model with covariates, we note that while Proposition 3 imposes more structural constraints such as a linear index, it offers a direct pathway to an estimation procedure that naturally accommodates the discreteness of the outcome, as we will outline in Section 7. This presents a clear estimation advantage compared to the methods proposed by Athey and Imbens (2006), which are not specifically tailored to discrete outcomes or the inclusion of covariates. However, it is important to stress that, akin to Theorem 4.4 in Athey and Imbens (2006), the identification result in Proposition 3 relies on strong assumptions that, while commonly invoked in the literature on semiparametric discrete outcome models, may sometimes be hard to justify.
5 Underreported Outcomes
This section addresses the scenario where outcomes may be underreported. It considers cases where, for $d\in\{0,1\}$, $C_{gt}(d)\neq Y_{gt}(d)$ with positive probability, a situation often referred to as “partial observability”. In nonlinear models, such discrepancies between true and reported outcomes generally lead to a failure of (point) identification. This issue persists even in fully parametric models unless additional exclusion restrictions are assumed (cf. Poirier, 1980).
Building upon the observation rule outlined in (1) and the latent outcome formulation of $R_{gt}(0)$ in (7), and drawing on the model for $Y_{gt}^{\ast}(0)$ presented in (13) from Section 4.2, we are now prepared to introduce a model for $R_{gt}^{\ast}(0)$. Similar in structure to the one for $Y_{gt}^{\ast}(0)$, this model incorporates additional covariates $\mathbf{Z}_{gt}$:
@@ 5 Underreported Outcomes
where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$. The covariate vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are allowed to share common elements. It is required, though, that each vector includes at least one distinct continuous variable not present in the other. W.l.o.g., we will assume that all elements, except for the first, in both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are common. Consequently, the augmented vectors $\overline{\mathbf{X}}_{gt}$ and $\overline{\mathbf{Z}}_{gt}$ are of the same dimension $k+1$.
@@
where $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$ denote CDFs of $-\varepsilon_{gt}$ and $-\nu_{gt}$, respectively, and $\overline{C}_{gt}(\cdot,\cdot)$ denotes a nonparametric Copula function. As noted in Section 2, the above setup bears similarities with the distributional sample selection model proposed by Chernozhukov et al. (2023), as well as the bivariate Probit model with a binary endogenous regressor considered by Han and Vytlacil (2017). However, a crucial distinction of our model, represented by (15), lies in the fact that the reporting intention $R_{gt}(0)$ is not directly observable. As a result, the second argument in the stated expression cannot be separately identified and estimated. This suggests that the identification results presented in the aforementioned papers are not directly applicable to our context. Specifically, the approach of using a single excluded variable in the second argument, as employed in Theorem 2.1 by Chernozhukov et al. (2023) and Theorem 6.1 by Han and Vytlacil (2017), may not be sufficient for point identifying the parameters inside the marginal distribution in our model, nor for the marginal distribution itself. This insight is supported by the simulation results in Section 8. These simulations suggest that, even within the parametric framework and using fairly large sample sizes, some parameters can still be poorly estimated if only one exclusion restriction is in place.
We make the following assumption:
@@
(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, the variables $X_{1,gt}$ and $Z_{1,gt}$ are excluded from the vectors $\mathbf{Z}_{gt}$ and $\mathbf{X}_{gt}$, respectively, while the other elements of them are common.
@@
with probability one, where $\mathbf{{W}}_{gt}$ denote the vector of regressors that are common to both $\mathbf{{X}}_{gt}$ and $\mathbf{{Z}}_{gt}$ and $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{W}_{gt}^{\prime}\right)^{\prime}$.
(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
Assumption 6(i) closely resembles Assumptions 1-2. While the covariate independence is restrictive, as previously discussed, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$. This allowance caters to scenarios where certain types of “consumers” may be more or less likely to misreport than others. This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000a). On the other hand, the other parts of Assumption 6 impose further conditions on the joint and marginal distributions of the unobservables. In particular, the real analytic property will allow for identification of the counterfactual distribution based on an extrapolation argument, which has also been featured in the microeconometric (see Arellano and Bonhomme, 2017; Pereda-Fernández, 2023) and the Industrial Organization identification literature (see Iaria and Wang, 2022, and references therein). Specifically, the extrapolation is based on the fact that if two real analytic functions coincide on an open neighborhood, then they coincide everywhere, and in particular also in the limit (if it exists). Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; D’Haultfoeuille and Maurel, 2013). Moreover, note that the real analytic property is satisfied by several well-known Copula families such as the Gaussian, Frank, Clayton, Plackett, Joe, or Gumbel Copula families. Indeed, while Assumptions 7(i) spells out the exclusion restrictions, Assumptions 7(ii) “only” imposes mild conditions on continuity of the excluded variables $X_{1,gt}$ and $Z_{1,gt}$, without requiring a “special regressor” as often necessitated in the semiparametric identification literature on discrete response models. On the other hand, the need for an excluded (continuous) variable in each decision equation is common to other identification work dealing with multiple latent outcome equations, see for instance Lewbel et al. (2021). Finally, Assumption 7(iii) extends Assumption 5 to the multivariate case.
We obtain the following identification result:
@@
Suppose Assumptions 6-7 hold and consider the models in (13)-(14). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\lambda_{gt}$, $\zeta_{gt}$, $F_{-\varepsilon_{gt}}(v)$ and $F_{-\nu_{gt}}(v)$ are identified. In particular, it holds for the counterfactual quantities:
@@
and

### v3

@@ 4.1 Partial Identification
Suppose $Y^{*}_{gt}(0)$ is generated by a CiC model, Assumptions 1-5 hold, and $\mathcal{Y}_{\mathbf{x},gt}=\{0,1,2\}$ for $(g,t)\neq(1,1)$. Suppose $\alpha\geq\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\neq(1,1)}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$. Then, for each $j\in\{0,1,2\}$ and $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$:
$F_{Y_{11}(1)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L_{11,\alpha}\left(j\mid\mathbf{x}\right),U_{11}\left(j\mid\mathbf{x}\right)\right],$
$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right),U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)\right],$
where
$L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=L_{10,\alpha}\left(\left(U_{00}\right)^{(-1)}\left(L_{01,\alpha}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
$U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=U_{10}\left(\left(L_{00,\alpha}\right)^{-1}\left(U_{01}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
and
$L_{gt,\alpha}\left(j\mid\mathbf{x}\right)$ $:=\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}}{1-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}},$ (5)
$U_{gt}\left(j\mid\mathbf{x}\right)$ $:=\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)$ (6)
for $j\in\{0,1\}$, and $L_{gt,\alpha}\left(j\mid\mathbf{x}\right)=U_{gt}\left(j\mid\mathbf{x}\right):=1$ for $j=2$. Moreover, the bounds are sharp when $\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}\leq\alpha<\min_{(g,t)\in\{0,1\}\times\{0,1\}}F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right).$
In addition, if $\alpha<\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$, then the set of admissible models is empty.
The above proposition derives valid bounds for the conditional cumulative distribution function of the counterfactual $Y_{11}(0)$ and the factual $Y_{11}(1)$. In addition, note that the bounds in Equation (5) and (6) are also valid bounds for the conditional distribution $Y(0)$ of each sub-group with $(g,t)\neq(1,1)$. In fact, when $\alpha=1$ and no bound is placed on underreporting, it is easy to see that
$\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\in\left[0,\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)\right]$
for $(g,t)\neq(1,1)$ and $j\in\{0,1\}$, which in turn implies that the bounds for the counterfactual distribution become
$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[0,1\right]$
for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. On the other hand, observe that Proposition 2 does also not rule out the case of no underreporting when the researcher sets $\alpha=0$. To see this, observe that when $\alpha=0$, it holds that:
$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=0$
for all $j\in\{0,1\}$. Under Assumption 3 and the first part of Assumption 4, we can write:
$F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)$
$=$ $\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)$
$\quad+\underbrace{\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)}_{=0}\times\left(1-\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\right).$
However, under Assumption 3 and $\alpha=0$, this equality holds if and only if $F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)=F_{C_{gt}\mid\mathbf{X}_{gt}}(j\mid\mathbf{x})$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and $j\in\{0,1\}$, which in turn restricts the set of admissible, observable distributions. [^6] Note that in this case $\alpha=\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}=0$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. Finally, in Section S.3.1 of the supplementary material, we comment on the possibility to estimate these bounds nonparametrically and to carry out inference via the nonparametric i.i.d. bootstrap following the ideas of Fang and Santos (2019) and Poirier and Masten (2020).
4.2 Point Identification
As the expressions for the bounds of the counterfactual distribution $Y_{11}(0)$ are only informative when paired with a relatively small upper bound on the cumulative underreporting probability (even the bounds may still be wide in practice), we take a different route in this subsection. Specifically, we introduce a semiparametric CiC model that jointly accounts for both consumption and reporting decisions. While this approach is more restrictive, as it imposes additional structure on the model, it will allow us to define and to identify new treatment parameters of interest, such as the DTT on the reporting intention.
To formalize the approach, let $R(d)$, $d\in\{0,1\}$ denote the potential “reporting intention”, analogous to $Y(d)$, with $R(d)$ taking on $J+1$ discrete value values. Here, we interpret $R(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully. The potential outcome of interest, $C(d)$, $d=0,1$, is then given by
$C(d)=\min\left\{Y(d),R(d)\right\}.$ (7)
Thus, respondents will truthfully report their consumption $Y(d)$ if and only if it is less than or equal to the reporting intention $R(d)$. For instance, if $Y(0)=2$ and $R(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2. It is worth noting that this formulation of underreporting behavior in (7) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019). The above definition also bears similarities with the binary selection setup in Klein et al. (2015). However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)). That is, unlike a selection setup where the selection indicator is fully observed, $R(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions. Importantly, also note that in our setup reporting behavior is allowed to change in response to the treatment status $d$. For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully.
The introduction of $R(d)$ allows us to define new policy parameter of interest such as the for the treated on the reporting intention:
$\tau^{r}(j\mid\mathbf{z}):=$ $\Pr\left(R(1)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right)-\Pr\left(R(0)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right),$ (8)
for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}$ given $G=1$ and $T=1$, where $\mathbf{Z}$ denotes a covariate vector nonnested within $\mathbf{X}$. The parameter in (8) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment. This may also shed light on how factors like stigma and shame influence the reporting behavior. In analogy to $Y^{\ast}(0)$, we conclude by defining the latent reporting intention variable $R^{\ast}(0)$:
$R(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}(0)\leq\iota_{3},\end{array}\right.$ (12)
4.2.1 Counterfactual Distribution
Building upon the observation rule outlined in (7) and the latent outcome formulation of $R(0)$ in (12), we consider semiparametric versions of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and further structure is placed on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively. Specifically, for every $(g,t)\in\{0,1\}$, we assume that $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}>0$. Thus:
@@ 4.2.1 Counterfactual Distribution
Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the mean of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter. Similar to the structure of $Y_{gt}^{\ast}(0)$, we also model $R_{gt}^{\ast}(0)$ as a function of additional covariates $\mathbf{Z}_{gt}$:
@@
where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$, $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$ and $\zeta_{gt}>0$. The covariate vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are allowed to share common elements. Depending on the treatment parameter of interest, however, it will be required that one vector includes a distinct continuous variable not present in the other.
@@
where $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$ denote CDFs of $-\varepsilon_{gt}$ and $-\nu_{gt}$, respectively, and $\overline{C}_{gt}(\cdot,\cdot)$ denotes a nonparametric Copula function. As noted in Section 2, the above setup bears similarities with the distributional sample selection model proposed by Chernozhukov et al. (2023), as well as the bivariate Probit model with a binary endogenous regressor considered by Han and Vytlacil (2017). However, a crucial distinction of our model, represented by (15), lies in the fact that the reporting intention $R_{gt}(0)$ is not directly observable. As a result, the second argument in the stated expression cannot be separately identified and estimated. This suggests that the identification results presented in the aforementioned papers are not directly applicable to our context. This insight is supported by the simulation results in Section LABEL:sec:simulation of the supplementary material. These simulations suggest that parameters from the model can, in some cases, still be poorly estimated even if one exclusion restriction is in place.
In what follows, let $X_{1,gt}$ ($Z_{1,gt}$) denote the first element of the vector $\mathbf{X}_{gt}$ ($\mathbf{Z}_{gt}$), and $\mathbf{X}_{-1,gt}$ ($\mathbf{Z}_{-1,gt}$) the vector containing the remaining ones. We make the following assumption:
@@
(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, (a) the variable $X_{1,gt}$ is excluded from the vector $\mathbf{Z}_{gt}$; (b) the variable $Z_{1,gt}$ is excluded from the vector $\mathbf{X}_{gt}$.
@@
with probability one, where $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{X}_{-1gt}^{\prime},\mathbf{W}_{-1gt}^{\prime}\right)^{\prime}$.
(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{d_{X_{gt}}+1}$ and $\mathbb{R}^{d_{Z_{gt}}+1}$, respectively, where $d_{X_{gt}}$ and $d_{Z_{gt}}$ denote the dimension of the vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$.
While the covariate independence in Assumption 6(i) is restrictive, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$. This allowance caters to scenarios where certain types of “consumers” may be more or less likely to misreport than others. This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000). On the other hand, the other parts of Assumption 6 impose further conditions on the joint and marginal distributions of the unobservables. In particular, the real analytic property will allow for identification of the counterfactual distribution based on an extrapolation argument, which has also been featured in the microeconometric (see Arellano and Bonhomme, 2017; Pereda-Fernández, 2023) and the Industrial Organization identification literature (see Iaria and Wang, 2022, and references therein). Specifically, the extrapolation is based on the fact that if two real analytic functions coincide on an open neighborhood, then they coincide everywhere, and in particular also in the limit (if it exists). Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; Corradi and Gutknecht, 2023). Moreover, note that the real analytic property is satisfied by several well-known Copula families such as the Gaussian, Frank, Clayton, Plackett, Joe, or Gumbel Copula families. Indeed, while Assumptions 7 (i)-(a) as well as (i)-(b) postulate an exclusion restriction for the identification of the parameters of the consumption and the reporting decision, respectively, part (ii) is rather “mild” in its continuity requirement. [^7] Note that Assumption 7(i)-(b) together with Assumption 6(i) imply Assumption 3 from Subsection 4.1. Specifically, it requires one continuous element in each vector, one of which is the instrumental variable. Finally, note that under the assumption that both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ contain an excluded, continuous variable, we may actually not only identify counterfactual parameters for either the consumption or the reporting decision as in Proposition 3 below, but for both simultaneously. We obtain the following identification result:
@@
Suppose Assumptions 6 holds and consider the models in (13)-(14). Also, suppose that the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as for $R_{gt}^{*}(0)$, and that $\mathcal{X}_{11}\subseteq_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
(i) Under Assumption 7 (i)-(a), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$, and $F_{-\varepsilon_{gt}}(v)$ are identified. In particular, we have that:
@@
(ii) Under Assumption 7 (i)-(b), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\zeta_{gt}$, and $F_{-\nu_{gt}}(v)$ are identified. In particular, we have that: